break it down for me..
3/4y + 2 = 1/4y +5

and also..
5(n+1) = 2(n+7)

Answer by stanbon(75887)

3/4y + 2 = 1/4y +5
(3/4 -1/4)y = 3
1/2 y = 3
y = 6
Cheers,
Stan H.